How would I write an equation of the line containing the point (3, -2) and perpendicular to the line y=3x-4. Show all steps. DO NOT GRAPH!

Answer by rfadrogane(214)

How would I write an equation of the line containing the point (3, -2) and perpendicular to the line y=3x-4.
Sol'n:
from the line equation:
y=mx+b
m-slope
m1=3
then the perpendicular slope is -1/m
m2=-1/3
at point (3,-2)
using point-slope form:
y-y1=m(x-x1)
y+2=-1/3(x-3)
3(y+2)=-(x-3)
3y+6=-x+3
x+3y+3=0 ----answer
^_^